### v2

# ReCode: Unify Plan and Action for Universal Granularity Control

Abstract
Real-world tasks require decisions at varying granularities, and humans excel at this by leveraging a unified cognitive representation where planning is fundamentally understood as a high-level form of action. However, current Large Language Model (LLM)-based agents lack this crucial capability to operate fluidly across decision granularities. This limitation stems from existing paradigms that enforce a rigid separation between high-level planning and low-level action, which impairs dynamic adaptability and limits generalization. We propose ReCode (Recursive Code Generation), a novel paradigm that addresses this limitation by unifying planning and action within a single code representation. In this representation, ReCode treats high-level plans as abstract placeholder functions, which the agent then recursively decomposes into finer-grained sub-functions until reaching primitive actions. This recursive approach dissolves the rigid boundary between plan and action, enabling the agent to dynamically control its decision granularity. Furthermore, the recursive structure inherently generates rich, multi-granularity training data, enabling models to learn hierarchical decision-making processes. Extensive experiments show ReCode significantly surpasses advanced baselines in inference performance and demonstrates exceptional data efficiency in training, validating our core insight that unifying planning and action through recursive code generation is a powerful and effective approach to achieving universal granularity control. The code is available at https://github.com/FoundationAgents/ReCode.
1 Introduction
Human decision-making in the real world naturally operates across varying granularities, seamlessly integrating high-level planning with fine-grained actions. Consider the simple act of preparing breakfast: one effortlessly shifts from deciding a high-level plan like “making bacon and eggs” to executing detailed, motor-level actions such as “cracking an egg”. Cognitive science attributes this fluid adaptability to the human brain’s integrated approach to control, which allows for decisions to be managed across different granularities. This integration is reflected in the brain’s shared representations for related cognitive processes (Prinz, 1997) and its hierarchical structure (Koechlin et al., 2003; Badre & D’esposito, 2009). Thus, humans naturally master this dynamic control of decision granularity, an essential capability for adaptive, intelligent behavior.
Achieving this adaptive intelligence is a primary goal for LLM-based agents (Liu et al., 2025), yet current frameworks fall short because they operate at fixed granularities. The core issue is that existing approaches explicitly separate planning from action into distinct decision-making processes. For example, ReAct (Yao et al., 2023) agent as shown in Figure 1(a), alternates strictly between reasoning and primitive actions step-by-step, limiting it to fine-grained decision-making without strategic foresight. On the other hand, agent with planner module (Figure 1(b)) separates high-level planning from low-level action using predefined structures. Such rigid boundaries impede agents from dynamically adjusting their decision granularity in response to evolving task complexities, ultimately resulting in brittle performance in complex, real-world environments.
To address this failure, our key insight is that planning and action are not fundamentally distinct cognitive processes, but rather represent decisions at different levels of granularity. At its essence, a plan is simply a higher-level action, analogous to how pseudo code represents higher-level conceptual thinking compared to concrete, executable code. Motivated by this insight, we propose ReCode (Recursive Code Generation), a novel paradigm that applies this idea by unifying plan and action within a single code representation. As illustrated in Figure 1(c), ReCode implements this pseudo-code-like concept by treating high-level plans as abstract placeholder functions, which the agent then recursively decomposes into finer-grained sub-functions until reaching executable primitive actions. This recursive process inherently dissolves the rigid boundary between planning and action, enabling the agent to dynamically control its decision granularity.
Furthermore, ReCode’s recursive structure provides a significant training advantage by inherently generating rich, hierarchical data. Unlike traditional approaches that yield flat action sequences, ReCode naturally produces structured decision trees capturing the entire cognitive process, from high-level planning down to executable actions. This comprehensive, multi-granularity training data significantly enhances the agent’s ability to learn complex task decompositions and adaptive decision-making strategies, improving both generalization and data efficiency.
Extensive experiments conducted across diverse, complex decision-making benchmarks validate the effectiveness of the ReCode paradigm. ReCode consistently surpasses strong paradigm baselines such as ReAct (Yao et al., 2023) and CodeAct (Wang et al., 2024b), delivering substantial improvements in overall task performance and problem-solving efficiency. Furthermore, our experiments reveal that the hierarchical, multi-granular data generated by ReCode plays a critical role in training flexible decision-making capabilities. These empirical results demonstrate ReCode’s potential as a transformative approach to building more adaptive, capable, and generalizable LLM-based agents.
In summary, our contributions are as follows:
We identify the lack of flexible decision granularity control as a fundamental limitation in existing LLM-based agents, which stems from the rigid separation of plan and action.
We propose ReCode, a novel agent paradigm that achieves universal granularity control by unifying plan and action within a single code representation. This unification is realized through a recursive mechanism where the agent decomposes high-level placeholder functions until primitive actions.
We demonstrate that this recursive structure inherently generates hierarchical, multi-granularity data, which enables models to learn hierarchical reasoning processes and leads to significant improvements in training data efficiency.
We validate our approach through comprehensive experiments, showing that ReCode significantly surpasses strong baselines in both inference performance and training efficiency across diverse benchmarks.
2 Related Work
The powerful reasoning and language capabilities of Large Language Models (LLMs) have inspired significant research effort to create agents that can interact with complex environments and accomplish specific tasks (Ahn et al., 2022; Huang et al., 2023; Yao et al., 2023). These efforts have largely fallen into two dominant paradigms: agents following the ReAct paradigm (Yao et al., 2023) and agents with explicit planners. However, both approaches, in their current forms, struggle with the fundamental limitation of a rigid separation between high-level planning and low-level action, which prevents flexible decision granularity control.
LLM-based ReAct Agent
Among these, paradigms like ReAct (Yao et al., 2023) are foundational, interleaving reasoning and primitive actions in a step-by-step loop. Code has also emerged as an expressive medium for actions, with paradigms like CodeAct (Wang et al., 2024b) leveraging the extensive code pre-training of LLMs. These foundational paradigms have spawned numerous agent systems (Yang et al., 2023a; Wu et al., 2024; Zhang et al., 2025; Liang et al., 2025; Hu et al., 2025). Much of the subsequent research has focused on optimizing this step-by-step loop, such as through synthetic data generation methods (Zelikman et al., 2022; Yang et al., 2024b; Wang et al., 2024a) and novel training algorithms (Song et al., 2024; Qiao et al., 2024; Du et al., 2024; Lee et al., 2024; Xia et al., 2025; Xiong et al., 2025; Cao et al., 2025). While these optimization efforts improve performance, they do not address the paradigm’s fundamental lack of integrated, high-level planning, which limits its foresight in complex, long-horizon tasks.
LLM-based Agent Planning
Planning emerged as a distinct research thrust, specifically to address this lack of foresight in ReAct Agent paradigm. Early approaches often adopted a “plan-and-execute” strategy, where a complete natural language plan is generated upfront before any actions are taken (Wang et al., 2023; Yang et al., 2023b; Kagaya et al., 2024; Sun et al., 2024a). However, these static plans are often brittle and difficult to adapt to dynamic environments. To address this, more sophisticated methods emerged, such as hierarchical planning frameworks (Paranjape et al., 2023; Sun et al., 2024b) and dynamic re-planning methods (Sun et al., 2023; Prasad et al., 2024), which continuously modify their plans based on environmental feedback. Nonetheless, these traditional planning methods still share the fundamental flaw of maintaining a rigid separation between the high-level plan and the low-level action execution.
While some recent approaches have begun to bridge this gap by explicitly integrating recursion or code-based processes (Liu et al., 2024; Schroeder et al., 2025; Zhang & Khattab, 2025), they still do not achieve a full unification and adaptive granularity control. It is highlighted that the core limitation across all existing paradigms remains this rigid separation of plan and action. In contrast, our work is built on the key insight that plan and action are not distinct processes, but rather represent decisions at different levels of granularity. This forms the foundation for our paradigm, which is unifying plan and action in a single representation.
3 Methodology
3.1 Preliminary
LLM-based Agent Decision Process
We model the interaction between an LLM-based agent and its environment as a simplified decision-making process , where is the state space, is the primitive action space, is the observation space, is the transition function, and is the reward function. At each step, the agent receives an observation and produces decisions that ultimately translate into executable primitive actions .
Beyond the primitive action space, we introduce the plan space , which encompasses intentions and goals at coarser granularities that cannot be directly executed but must be refined into sequences of primitive actions or intermediate sub-plans. We define the decision space as , representing all possible outputs an agent can generate across different granularities.
Current agent paradigms typically operate with predefined, fixed decision spaces. ReAct agents directly select from a pre-determined action set , while planner-based agents generate sequences over the same fixed or rely on manually specified plan templates. This constraint fundamentally limits adaptability, as agents cannot dynamically generate novel decisions suited to unforeseen contexts.
Decision Granularity
Real-world tasks demand decisions at varying granularities. Fine-grained decisions in correspond to immediately executable primitive actions such as run('crack egg'), while coarse-grained decisions in represent higher-level intentions requiring decomposition, such as prepare_breakfast().
The granularity forms a natural hierarchy. Consider the breakfast preparation example. The decision “prepare breakfast” is coarser than “cook eggs”, which is in turn coarser than “crack egg”. Each level encompasses broader objectives and longer temporal horizons. The finest level consists of executable primitive actions, while the coarsest level represents the complete task specification.
Motivated by the perspective that plans are essentially high-level actions at different abstraction levels, ReCode unifies decisions across all granularities within a single code representation. High-level plans in take the form of placeholder functions, which are recursively refined into finer-grained components until reaching executable primitive actions in . This formulation enables the agent to dynamically generate decisions, whether plans or actions, at appropriate granularities for the given context, effectively creating an unbounded decision space.
3.2 Method Overview
We introduce ReCode, an LLM-based agent paradigm with recursive code generation that achieves universal control of decision granularity. This paradigm begins by unifying plan and action into the same representation, enabling the entire decision-making process to be implemented through a single, consistent operation. This unification naturally gives rise to a tree structure where the highest-level task can be recursively decomposed until executable leaves. This dynamic process of building and executing a decision tree is illustrated in Figure 2.
Unify Plan and Action
The key insight of ReCode is that plans and actions, despite their apparent differences, can be unified under a single executable code representation. This unification addresses a fundamental limitation in current LLM-based agent frameworks: the rigid separation between abstract planning and concrete primitive action, which prevents flexible decision granularity control.
We represent both plans and actions as Python function calls. This establishes a common computational substrate for the agent’s policy. Actions are executable and environment-specified that can directly interface with the environment, such as run('click the submit button'). Plans are expressed as unimplemented placeholder functions. These functions encapsulate sub-goals that require further expansion, like prepare_breakfast() and get_ingredients(). For example, a high-level plan prepare_breakfast() can be expanded by the policy into a code block containing other plans, like get_ingredients() and cook_meal(). These, in turn, are recursively broken down until they resolve into a sequence of primitive actions like run('open refrigerator') and run('turn on stove').
Recursive Code Generation
Building on this unified representation, ReCode operates via a recursive generation and execution loop, as detailed in Algorithm 1. The process begins with a rule-based text-to-code transformation that converts the task instruction and initial observation into a placeholder function, which is the root node. The agent’s policy model then expands the current node into a child code block.
An executor processes the generated code block sequentially. When encountering a primitive action, the executor directly executes it in the environment. When encountering a placeholder function, the executor triggers an expansion by invoking ReCode again with this placeholder as the new current node. The decision tree grows in this process, gradually breaking down the task into more detailed plans until all leaves are executable primitive actions.
The recursion terminates when a generated code block contains only primitive actions. Execution then returns to the previous, coarser level. This process continues until the initial top-level placeholder is fully resolved, allowing the agent to seamlessly navigate across granularity levels. The policy adaptively determines when to maintain abstract planning and when to commit to concrete actions based solely on the current decision context, without explicit granularity supervision. The resulting execution trace forms a hierarchical decision tree that captures the complete reasoning process from strategic planning to primitive execution.
3.3 Implementation Details
Bridging the ReCode paradigm from theory to practice requires several key engineering efforts to ensure robust and stable agent performance.
Task Initialization
The starting point for ReCode is converting the natural language task instruction into the root placeholder function. We employ a simple rule-based method that directly encapsulates the task instruction and initial observation as string arguments within a predefined template, such as solve(instruction, observation). This task-agnostic approach delegates the full responsibility of interpreting and decomposing the task to the learned policy model, compelling it to develop core planning capabilities from raw inputs and enhancing generalization across diverse tasks and environments.
Context Management
ReCode implements context management through a unified variable namespace that persists throughout task execution. When expanding a placeholder function, the system serializes all currently provided variables and their values into a structured text description, which is injected into the prompt as context. After executing the generated code block, newly created or modified variables are updated in this shared namespace. This design creates a hierarchical information flow where sub-tasks can access context established by parent tasks at any level of the call stack. Crucially, the policy model only sees the current variable state, not the full execution history. This enforces explicit state management. The model must learn to consciously save important information to variables for future use. For example, capturing an action’s output as obs = run('go to cabinet 1') allows subsequent inspection of the obs variable. This approach maintains concise and relevant context while guiding the model toward structured and deliberate planning.
Error Handling and Recursion Control
To ensure robust execution, ReCode addresses two practical challenges. First, code generated by LLMs is susceptible to syntax or runtime errors. We employ a self-correction loop during inference. Upon execution failure, the system re-invokes the policy with the original placeholder and the error traceback as additional context, allowing it to recover from transient generation failures within a single trajectory. Second, unbounded recursion can lead to infinite loops or excessively deep decomposition trees. We impose a maximum recursion depth of 10 for our benchmark tasks, chosen as a conservative upper bound above the empirically optimal depth (in Section 4.2), balancing planning complexity with guaranteed termination.
4 Experiment
4.1 Experimental Setup
Environments
We conduct experiments across three text-simulated environments that represent diverse decision-making challenges for LLM-based agents. ALFWorld (Shridhar et al., 2021) provides an embodied household environment where agents navigate and manipulate objects to complete daily tasks. WebShop (Yao et al., 2022) simulates an online shopping scenario requiring agents to search, compare, and purchase products based on specific requirements. ScienceWorld (Wang et al., 2022) presents a scientific laboratory setting where agents conduct experiments and manipulate scientific instruments. All three environments are partially observable Markov decision processes where agents cannot fully infer the environment state from observations alone. For more detailed descriptions of the environments, please refer to Appendix A.1.
Baselines
To evaluate the effectiveness of ReCode, we divide our experiments into the inference part and the training part.
For inference experiments, we compare against several mainstream paradigms: ReAct (Yao et al., 2023), which alternates between reasoning and action for iterative environment interaction; CodeAct (Wang et al., 2024b), extending ReAct’s action space from natural language to Python code. And some of the work focused on improving LLM-based agent planning: AdaPlanner (Sun et al., 2023), which pre-writes action sequences in Python and iteratively modifies them based on environmental feedback; and ADaPT (Prasad et al., 2024), which decomposes tasks on-demand using ReAct as the sub-task executor.
| Method | Metric | ALFWorld | ScienceWorld | WebShop |
| ReAct | Data Pairs | 27,607 | 12,833 | 7,181 |
| Tokens | 18,958,782 | 15,356,801 | 9,409,144 | |
| Avg. Reward% (Std) | 100.0∗ | 100.0 0.00 | 82.3 21.15 | |
| CodeAct | Data Pairs | 5,049 | 5,984 | 3,853 |
| Tokens | 9,630,822 | 15,189,063 | 9,214,185 | |
| Avg. Reward% (Std) | 100.0∗ | 60.8 24.51 | 94.6 10.36 | |
| ReCode | Data Pairs | 6,385 | 3,500 | 2,473 |
| Tokens | 4,962,898 | 3,986,769 | 2,662,276 | |
| Avg. Reward% (Std) | 100.0∗ | 88.5 14.58 | 97.5 7.05 |
For training experiments, which evaluate how well a paradigm’s data structure lends itself to learning, we conduct supervised fine-tuning (SFT) (Wei et al., 2022) experiments on ReCode, ReAct, and CodeAct to directly compare their learning efficiency. Also, we compare some advanced training methods on the ReAct-style agent, including ETO (Song et al., 2024), which learns from expert trajectories via Direct Preference Optimization (Rafailov et al., 2023); and WKM (Qiao et al., 2024), introducing parameterized world models for training on expert and synthetic trajectories. Due to differences in models, training setups, or environment configurations, we could not directly implement these latter works, but list them for reference.
Evaluation
We report average reward (%) as our primary metric across all experiments. The environments employ different reward settings. ALFWorld provides binary rewards (0 or 1) indicating task completion, while WebShop and ScienceWorld offer dense rewards ranging from 0 to 1 based on task progress and accuracy. ALFWorld and ScienceWorld both include out-of-distribution evaluation sets to assess generalization capabilities.
Implementation Details
For inference experiments, all methods use GPT-4o mini (Hurst et al., 2024) to ensure fair comparison. We adapt few-shot prompts from prior work, making minimal modifications to fit each method’s format while keeping the number of examples consistent. The few-shot examples used for ReCode are provided in Appendix A.3.
For training experiments, we use Qwen2.5-7B-Instruct (Yang et al., 2024a) as the base model. We construct the training datasets through the following process: First, for each training instance, we use a powerful teacher model, DeepSeek-V3.1 (DeepSeek-AI, 2024), to generate trajectories using ReAct, CodeAct, and ReCode respectively. Second, we filter these trajectories by keeping only the top 40% based on final reward. Third, from each retained trajectory, we extract input-output pairs as SFT training data, where the input excludes few-shot examples to focus on the model’s learning of the paradigm itself.
Table 1 presents detailed statistics of the resulting training datasets. As shown, different paradigms produce vastly different amounts of training data from the same pool of source instances. ReCode generates significantly fewer but more compact data pairs (e.g., 6,385 pairs with 4.96M tokens on ALFWorld) compared to ReAct (27,607 pairs with 18.96M tokens). All models are trained using full-parameter fine-tuning on 8 NVIDIA A800 80GB GPUs with the open-source verl framework (Sheng et al., 2024).
4.2 Inference Results
| Method | ALFWorld | ScienceWorld | WebShop | Average | ||
| Seen | Unseen | Seen | Unseen | |||
| ReAct GPT-4o mini | 59.29 | 64.18 | 47.02 | 38.91 | 27.62 | 47.4 |
| CodeAct GPT-4o mini | 72.14 | 85.07 | 22.68 | 18.41 | 21.37 | 43.9 |
| AdaPlanner GPT-4o mini | 75.00 | 92.54 | 28.49 | 26.11 | 29.17 | 50.3 |
| ADaPT GPT-4o mini | 34.29 | 41.04 | 37.53 | 32.94 | 32.79 | 35.7 |
| ReCode GPT-4o mini | 83.57 | 96.27 | 42.78 | 41.18 | 39.97 | 60.8 |
| ReAct Gemini 2.5 Flash | 85.71 | 85.07 | 24.73 | 25.60 | 39.74 | 52.2 |
| CodeAct Gemini 2.5 Flash | 52.86 | 63.43 | 23.89 | 20.64 | 43.67 | 40.9 |
| ReCode Gemini 2.5 Flash | 90.00 | 96.27 | 54.35 | 43.77 | 46.57 | 66.2 |
| ReAct DeepSeek-V3.1 | 90.71 | 88.06 | 68.52 | 61.75 | 22.94 | 66.4 |
| CodeAct DeepSeek-V3.1 | 51.43 | 59.70 | 35.96 | 27.91 | 36.50 | 42.3 |
| ReCode DeepSeek-V3.1 | 95.71 | 94.78 | 50.20 | 50.21 | 55.07 | 69.2 |
Main Results
As shown in Table 2, ReCode achieved significant performance improvements across all three environments, with an average score of 60.8, surpassing the best baseline method by 10.5 (relative 20.9%). Notably, compared to the ReAct and CodeAct paradigms, ReCode improved average performance by 13.4 and 16.9, respectively. The results demonstrate ReCode’s exceptional capabilities across diverse task domains. In the ALFWorld environment, our method exhibits particularly strong performance, achieving a high score of 83.57 on seen tasks and an outstanding 96.27 on unseen tasks, substantially outperforming all baseline methods. On WebShop, ReCode maintains competitive performance with a score of 39.97, representing a 21.9% improvement over ADaPT. In the ScienceWorld environment, while ReCode shows comparable performance to ReAct on seen data and notable improvement on unseen tasks. These consistent improvements across varied environments underscore the effectiveness of our approach in enhancing both task-specific performance and cross-domain adaptability. We provide detailed case studies in Section 4.4, showcasing representative trajectories from ALFWorld to illustrate ReCode’s decision-making process.
Multiple Model Results
To illustrate that the ReCode paradigm is applicable to most models, we also experiment with Gemini 2.5 Flash (Comanici et al., 2025) and DeepSeek-V3.1, with results shown in Table 2. The cross-model evaluation demonstrates ReCode’s robustness and generalizability beyond a single LLM. On Gemini 2.5 Flash, ReCode achieved an average performance of 66.2, maintaining consistent improvements over baseline methods. Similarly, when applied to DeepSeek-V3.1, it obtained a 69.2 average score, further validating the paradigm’s effectiveness. These results indicate that ReCode’s performance gains are not dependent on specific model, but rather stem from the fundamental improvements in the reasoning and code generation process.
Recursion Depth Analysis
To validate our depth control mechanism, we conduct ablation experiments with varying maximum depth limits on the ScienceWorld seen set using GPT-4o mini. As shown in Figure 3, performance exhibits a clear inverted-U pattern, peaking at depth 8 (43.22%). Shallow depths restrict hierarchical decomposition, while excessive depths decline performance, likely due to over-decomposition fragmenting the decision-making process. The optimal range lies between depths 6-12. Based on these findings, we set the maximum depth to 10 across all environments, providing a conservative upper bound that prevents unbounded recursion while rarely constraining practical reasoning.
Cost Analysis
As shown in Table 3, ReCode is remarkably cost-efficient. On average, a ReCode trajectory costs 78.9% less than ReAct and 84.4% less than CodeAct. This dramatic improvement stems from ReCode’s structured exploration. Instead of the extensive, token-intensive, step-by-step reasoning cycles characteristic of reactive methods, ReCode’s hierarchical decomposition allows the agent to make more abstract and potent decisions. This leads to shorter, more direct reasoning paths, significantly reducing the number of API calls and total tokens required to solve a task while achieving superior performance.
| Method | ALFWorld | ScienceWorld | WebShop | Average | ||
| Seen | Unseen | Seen | Unseen | |||
| ReAct | 10.55 | 9.76 | 10.29 | 11.07 | 3.15 | 8.96 |
| CodeAct | 5.24 | 5.92 | 13.42 | 12.13 | 24.04 | 12.15 |
| ReCode | 2.11 | 1.99 | 1.91 | 2.23 | 1.19 | 1.89 |
4.3 Training Results
This section demonstrates the effectiveness of ReCode’s hierarchical data structure for agent training. We present comprehensive results across three environments and analyze the relationship between data and performance.
| Method | ALFWorld | ScienceWorld | WebShop | Average | ||
| Seen | Unseen | Seen | Unseen | |||
| ReAct Qwen2.5-7B-Instruct | 57.86 | 62.69 | 19.49 | 14.85 | 32.33 | 37.4 |
| CodeAct Qwen2.5-7B-Instruct | 67.86 | 69.40 | 13.47 | 9.29 | 37.43 | 39.5 |
| ReAct+SFT Qwen2.5-7B-Instruct | 90.00 | 90.30 | 60.28 | 54.64 | 42.61 | 67.6 |
| CodeAct+SFT Qwen2.5-7B-Instruct | 86.43 | 88.06 | 37.25 | 33.59 | 33.48 | 55.8 |
| ReAct+ETO∗ Llama-3-8B-Instruct | 64.29 | 64.18 | 57.90 | 52.33 | 64.57 | 60.7 |
| ReAct+WKM∗ Llama-3-8B-Instruct | 68.57 | 65.93 | 60.12 | 54.75 | 66.64 | 63.2 |
| ReCode Qwen2.5-7B-Instruct | 66.43 | 73.88 | 15.84 | 15.38 | 42.24 | 42.8 |
| ReCode+SFT Qwen2.5-7B-Instruct | 92.14 | 97.01 | 59.85 | 52.39 | 50.85 | 70.4 |
Main Results
Table 4 presents the training results using Qwen2.5-7B-Instruct. ReCode+SFT achieves a strong average performance of 70.4% across all environments, surpassing both ReAct+SFT (67.6%) and CodeAct+SFT (55.8%).
Notably, this performance is achieved with remarkable data efficiency. As shown in Table 1, ReCode’s training set contains only 3,500 pairs, 3.7 times less than the 12,833 pairs used by ReAct, and its average reward (88.5%) is also lower than ReAct’s (100.0%). This demonstrates that ReCode+SFT can achieve superior results even with significantly less and potentially lower-quality training data. On individual benchmarks, ReCode+SFT shows dominant performance on ALFWorld and WebShop, and remains competitive on ScienceWorld.
Data Efficiency Analysis
To systematically investigate the source of ReCode’s superior data efficiency, we conducted training on data subsets filtered by reward percentile . Figure 4 and Table 5 present the results on ScienceWorld. The results reveal distinct scaling behaviors. ReCode achieves 55.34% on seen tasks at using only 1,713 pairs, improving to 59.85% at with 3,500 pairs. In contrast, ReAct requires 12,833 pairs at to reach 60.28% on seen tasks. At matched percentiles, ReCode consistently outperforms ReAct with substantially fewer data pairs. At , ReCode achieves 44.87% with 688 pairs versus ReAct’s 34.05% with 3,094 pairs (4.5 more). At , ReCode reaches 55.34% with 1,713 pairs versus ReAct’s 36.34% with 6,320 pairs (3.7 more). Figure 4 shows that ReCode’s curves rise steeply then plateau, while ReAct’s curves show gradual improvement requiring substantially more data, confirming that ReCode’s hierarchical structure provides richer learning signals per training example.
Figure 4: Performance of ReCode (blue) and ReAct (red) on ScienceWorld under different filtering percentiles. Performance on seen tasks is represented by triangles, and performance on unseen tasks is represented by squares.
Method
Percentile
Data Pairs
ScienceWorld
Seen
Unseen
ReAct
10
3,094
34.05
31.66
20
6,320
36.34
30.96
30
9,488
47.52
44.82
40
12,833
60.28
54.64
ReCode
10
688
44.87
38.33
20
1,713
55.34
49.93
30
2,499
57.63
50.97
40
3,500
59.85
52.39
Table 5: Training data statistics and corresponding ScienceWorld performance for ReAct and ReCode. The table lists the number of data pairs and the resulting average rewards for each method at different filtering percentiles ().
4.4 Case Study
To provide a concrete illustration of ReCode’s decision-making process, we analyze a representative trajectory from the ALFWorld environment for the task “put two alarmclock in dresser”. The complete execution trace and a diagram of the code generation flow are available in Figure 5 of the Appendix A.4.
Initially, the agent formulates hybrid decisions that integrate abstract sub-task placeholder functions with concrete, primitive actions. These initial decisions, shown in the following code block, include functions like find_and_take for complex procedures that require further decomposition, and known, simple steps like run('go to dresser 1') as primitive actions.
The agent then executes this hybrid code block sequentially. Primitive actions like run(...) are executed directly by the environment executor. While, placeholder functions such as find_and_take trigger an on-demand expansion, where the agent generates the necessary code to complete that specific sub-task just in time. This hybrid strategy allows the agent to maintain a high-level, strategic view of the task while committing to concrete actions only when the steps are simple and certain. It combines the robustness of hierarchical planning with the efficiency of direct execution, providing a flexible and powerful problem-solving framework that can adapt its level of abstraction as needed.
5 Conclusion
In this work, we introduced ReCode, a novel paradigm for LLM-based agents that achieves universal control of decision granularity through recursive code generation. By unifying plan and action within a single code representation, ReCode addresses fundamental limitations in current agent paradigms that rigidly separate decisions of different granularity. Our key insight is that planning is fundamentally high-level action at different granularities. This enables agents to dynamically decompose goals into hierarchical code structures without reliance on predefined action spaces. Extensive experiments across three diverse environments demonstrate ReCode’s dual advantages. In inference, ReCode achieves superior performance with over 20.9% improvement. In training, ReCode exhibits remarkable efficiency using dramatically fewer data while achieving better results. The recursive structure naturally generates rich, multi-granularity training data that captures complete cognitive processes from strategic planning to concrete action. ReCode establishes a foundation for more adaptive AI agents that can fluidly transition across decision granularities based on task complexity and situational demands.
6 Limitations and Future Work
While ReCode demonstrates significant potential, its effectiveness depends critically on the base model’s reasoning capabilities and few-shot example quality. The paradigm requires models to generate structured code within a specific framework, and any deficiency in planning ability compromises subsequent decision quality. As observed in our experiments, ReCode’s performance can be inconsistent, largely due to models’ difficulty adhering to the required code generation format or fully grasping the framework’s underlying logic.
These limitations point to clear research directions centered on agent learning. A primary focus should be enhancing models’ intrinsic ability to understand and operate within the ReCode framework through specialized pre-training objectives or comprehensive fine-tuning strategies. Reinforcement learning approaches could reward agents for creating efficient hierarchical plans, enabling them to optimize their own expansion processes. Additionally, exploring methods to make the framework more robust to generation errors and investigating automatic curriculum learning for progressive skill acquisition represent promising avenues. These directions would represent significant steps toward more autonomous and capable agent learning systems.
Appendix A Appendix
A.1 Datasets
ALFWorld
ALFWorld (Shridhar et al., 2021) provides a text-only game environment where the agent is required to complete a series of long-horizon household tasks based on natural language instructions. Built upon the embodied vision benchmark ALFRED (Shridhar et al., 2020), ALFWorld inherits its six core task types: pick and place, examine in light, clean and place, heat and place, cool and place, and pick two and place. After executing a command (e.g., “go to countertop 1”), the agent receives immediate textual feedback from the environment (e.g., “You arrive at countertop 1. On the countertop 1, you see an egg 3, a peppershaker 1, and a tomato 1”). ALFWorld only provides a binary reward (1/0) to indicate whether the task was successfully completed.
WebShop
WebShop Yao et al. (2022) is an online shopping website environment constructed with 1.18 million real-world product data points. The agent must navigate the site by searching and clicking buttons to browse various products, with the goal of finding and purchasing a specific item that meets the given requirements. We follow the data partitioning from MPO (Xiong et al., 2025), using a test set of 200 instances. WebShop provides a dense reward between 0 and 1 based on the agent’s actions and interactions with products. A task is considered successful only if the reward is strictly equal to 1.
ScienceWorld
ScienceWorld is a text-based simulation environment centered on completing basic scientific experiments. The agent needs to interact with the environment to measure certain values or to complete an experiment by applying scientific knowledge. The tasks in ScienceWorld are meticulously categorized, with each task having several variations and offering numerous different sub-goals. We adopt the data split from ETO (Song et al., 2024) and exclude Task-9 and Task-10 from the trajectories. ScienceWorld provides a dense reward between 0 and 1 for each task, based on the completion of its sub-goals.
The data statistics for each environment are summarized in Table 6.
| Train | Seen | Unseen | |
| ALFWorld | 3553 | 140 | 134 |
| WebShop | 1824 | 200 | - |
| ScienceWorld | 1483 | 194 | 211 |
A.2 Prompts
This appendix provides detailed prompt templates for ReAct, CodeAct and ReCode agents.
ReAct Agent
The ReAct agent employs a reasoning and acting cycle pattern, solving problems by alternating between thinking and acting. The following are the prompt templates used in ALFWorld, ScienceWorld, and WebShop.
CodeAct Agent
CodeAct agent solves problems by generating and executing Python code, interacting with the environment using specific functions.
ReCode Agent
ReCode Agent shares a set of prompt templates across all environments, providing different primitive action descriptions and examples based on the environment.
A.3 Fewshot Examples
We provide detailed examples used by the ReCode agent across different environments. Each example demonstrates the hierarchical decomposition approach where high-level tasks are progressively expanded into executable primitive actions.
ALFWorld Examples
The ReCode agent uses task-specific examples for different categories in ALFWorld. Each category represents a distinct household task pattern with slight variations in logic while maintaining consistent structure. Here we present the cool task category as a representative example.
ScienceWorld Examples
The ScienceWorld examples demonstrate scientific experiment tasks that require domain knowledge application and spatial reasoning. The agent must navigate between rooms, manipulate objects, and perform chemistry or physics operations. Here we present a representative chemistry task: creating green paint through color mixing.
A.3.1 WebShop Examples
The WebShop examples demonstrate e-commerce navigation tasks where the agent must search for products, evaluate options against specified requirements, and complete purchases. Note that since clickable buttons in the environment are represented with square brackets (e.g., [Search], [Buy Now]), we use INPUT: and OUTPUT: instead of [Input] and [Output] in these examples to avoid confusion.
A.4 Case Study
To provide a concrete illustration of ReCode’s decision-making process, we analyze a representative trajectory from the ALFWorld environment for the task “put two alarmclock in dresser”. The agent’s complete execution flow, derived from the log file, is detailed below. Figure 5 provides a high-level visualization of this recursive process.
The agent’s primary goal is to place two alarm clocks in a dresser. Given that it can only carry one object at a time, a sequential find-and-place strategy is necessary.
The following steps detail the agent’s turn-by-turn execution. It is important to note that this is not a pre-generated script. Instead, the agent starts with the high-level plan from Step 1 and executes it line by line. When the executor encounters an unimplemented placeholder function (e.g., declare_init_vars), execution pauses, and a recursive call is made to the LLM to generate the necessary code for that specific function. The newly generated code is then executed, and this cycle repeats until the entire plan is resolved.
Step 1: Initial High-Level Plan
First, the agent formulates an initial high-level plan by expanding the root solve function. This plan breaks down the complex task into a sequence of abstract placeholder functions without specifying low-level primitive actions. This strategic scaffold includes steps for finding and placing each of the two alarm clocks, as well as updating its internal knowledge about object locations.
Step 2: Declaring Initial Variables
The agent begins by executing the first placeholder, declare_init_vars. It expands this function to parse the initial instruction and observation, identifying key variables. This step grounds the agent by establishing the target object (“alarmclock”), the destination (“dresser 1”), and a comprehensive list of all searchable locations.
Step 3: Finding and Taking the First Alarm Clock
Upon successfully executing the variable declarations, the executor proceeds to the next line in the plan: the find_and_take placeholder. This triggers another on-demand expansion cycle, prompting the agent to generate the following code for a methodical search:
Step 4: Placing the First Alarm Clock
After the code from the previous step successfully executes and finds "alarmclock 3", the agent continues to the next task in its high-level plan, put_in. This function is then expanded to generate the code for navigating to the dresser and placing the object inside.
Step 5: Updating Location Knowledge
After placing the first clock, the agent must update its search space for the second clock. It expands update_all_location_IDs, which intelligently slices the location list to resume searching from the last successful spot (“desk 1”) and removes the destination (“dresser 1”) to prevent redundant checks.
Step 6 & 7: Finding and Placing the Second Alarm Clock
Finally, the agent repeats the find-and-place cycle by expanding find_and_take_again and put_in_again. It reuses the same logic as before but operates on the updated location list. It successfully finds "alarmclock 2" on "desk 1", picks it up, navigates back to the dresser, and places it inside, completing the task.

### v3

@@
 Cognitive science attributes this fluid adaptability to the human brain’s integrated approach to control, which allows for decisions to be managed across different granularities.
-This integration is reflected in the brain’s shared representations for related cognitive processes (Prinz, 1997) and its hierarchical structure (Koechlin et al., 2003; Badre & D’esposito, 2009).
+This integration is reflected in the brain’s shared representations for related cognitive processes (prinz1997perception) and its hierarchical structure (koechlin2003architecture; badre2009rostro).
 Thus, humans naturally master this dynamic control of decision granularity, an essential capability for adaptive, intelligent behavior.
-Achieving this adaptive intelligence is a primary goal for LLM-based agents (Liu et al., 2025), yet current frameworks fall short because they operate at fixed granularities.
+Achieving this adaptive intelligence is a primary goal for LLM-based agents (liu2025advances), yet current frameworks fall short because they operate at fixed granularities.
 The core issue is that existing approaches explicitly separate planning from action into distinct decision-making processes.
-For example, ReAct (Yao et al., 2023) agent as shown in Figure 1(a), alternates strictly between reasoning and primitive actions step-by-step, limiting it to fine-grained decision-making without strategic foresight.
+For example, ReAct (yao2023react) agent as shown in Figure 1(a), alternates strictly between reasoning and primitive actions step-by-step, limiting it to fine-grained decision-making without strategic foresight.
 On the other hand, agent with planner module (Figure 1(b)) separates high-level planning from low-level action using predefined structures.
@@
 Extensive experiments conducted across diverse, complex decision-making benchmarks validate the effectiveness of the ReCode paradigm.
-ReCode consistently surpasses strong paradigm baselines such as ReAct (Yao et al., 2023) and CodeAct (Wang et al., 2024b), delivering substantial improvements in overall task performance and problem-solving efficiency.
+ReCode consistently surpasses strong paradigm baselines such as ReAct (yao2023react) and CodeAct (wang2024executable), delivering substantial improvements in overall task performance and problem-solving efficiency.
 Furthermore, our experiments reveal that the hierarchical, multi-granular data generated by ReCode plays a critical role in training flexible decision-making capabilities.
@@
 2 Related Work
-The powerful reasoning and language capabilities of Large Language Models (LLMs) have inspired significant research effort to create agents that can interact with complex environments and accomplish specific tasks (Ahn et al., 2022; Huang et al., 2023; Yao et al., 2023).
-These efforts have largely fallen into two dominant paradigms: agents following the ReAct paradigm (Yao et al., 2023) and agents with explicit planners.
+The powerful reasoning and language capabilities of Large Language Models (LLMs) have inspired significant research effort to create agents that can interact with complex environments and accomplish specific tasks (ahn2022can; huang2023inner; yao2023react).
+These efforts have largely fallen into two dominant paradigms: agents following the ReAct paradigm (yao2023react) and agents with explicit planners.
 However, both approaches, in their current forms, struggle with the fundamental limitation of a rigid separation between high-level planning and low-level action, which prevents flexible decision granularity control.
 LLM-based ReAct Agent
-Among these, paradigms like ReAct (Yao et al., 2023) are foundational, interleaving reasoning and primitive actions in a step-by-step loop.
-Code has also emerged as an expressive medium for actions, with paradigms like CodeAct (Wang et al., 2024b) leveraging the extensive code pre-training of LLMs.
-These foundational paradigms have spawned numerous agent systems (Yang et al., 2023a; Wu et al., 2024; Zhang et al., 2025; Liang et al., 2025; Hu et al., 2025).
-Much of the subsequent research has focused on optimizing this step-by-step loop, such as through synthetic data generation methods (Zelikman et al., 2022; Yang et al., 2024b; Wang et al., 2024a) and novel training algorithms (Song et al., 2024; Qiao et al., 2024; Du et al., 2024; Lee et al., 2024; Xia et al., 2025; Xiong et al., 2025; Cao et al., 2025).
+Among these, paradigms like ReAct (yao2023react) are foundational, interleaving reasoning and primitive actions in a step-by-step loop.
+Code has also emerged as an expressive medium for actions, with paradigms like CodeAct (wang2024executable) leveraging the extensive code pre-training of LLMs.
+These foundational paradigms have spawned numerous agent systems (yang2023auto; wu2024autogen; zhang2025aflow; openmanus2025; hu2025owl).
+Much of the subsequent research has focused on optimizing this step-by-step loop, such as through synthetic data generation methods (zelikman2022star; yang2024react; wang2024learning) and novel training algorithms (song2024trial; qiao2024agent; du2024improving; lee2024rlaif; xia2025sand; xiong2025mpo; cao2025pgpo).
 While these optimization efforts improve performance, they do not address the paradigm’s fundamental lack of integrated, high-level planning, which limits its foresight in complex, long-horizon tasks.
@@
 Planning emerged as a distinct research thrust, specifically to address this lack of foresight in ReAct Agent paradigm.
-Early approaches often adopted a “plan-and-execute” strategy, where a complete natural language plan is generated upfront before any actions are taken (Wang et al., 2023; Yang et al., 2023b; Kagaya et al., 2024; Sun et al., 2024a).
+Early approaches often adopted a “plan-and-execute” strategy, where a complete natural language plan is generated upfront before any actions are taken (wang2023plan; yang2023intercode; kagaya2024rap; sun2024corex).
 However, these static plans are often brittle and difficult to adapt to dynamic environments.
-To address this, more sophisticated methods emerged, such as hierarchical planning frameworks (Paranjape et al., 2023; Sun et al., 2024b) and dynamic re-planning methods (Sun et al., 2023; Prasad et al., 2024), which continuously modify their plans based on environmental feedback.
+To address this, more sophisticated methods emerged, such as hierarchical planning frameworks (paranjape2023art; sun2024pearl) and dynamic re-planning methods (sun2023adaplanner; prasad2024adapt), which continuously modify their plans based on environmental feedback.
 Nonetheless, these traditional planning methods still share the fundamental flaw of maintaining a rigid separation between the high-level plan and the low-level action execution.
-While some recent approaches have begun to bridge this gap by explicitly integrating recursion or code-based processes (Liu et al., 2024; Schroeder et al., 2025; Zhang & Khattab, 2025), they still do not achieve a full unification and adaptive granularity control.
+LLM-based Agent with Recursive Structure
+While some recent approaches have begun to bridge this gap by explicitly integrating recursion or code-based processes, they still do not achieve a full unification and adaptive granularity control.
+Code-as-Policies (liang2023code) and follow-up methods (wang2023demo2code; mu2024robocodex; ahn2025towards) use complete executable programs to represent the agent policy, and recursively expand high-level sketches into full programs to allocate context across different sub-tasks, but lack mechanisms to dynamically adjust the plan based on interaction with the environment.
+THREAD and REPL-Plan (schroeder2025thread; liu2024interactive) instead introduce recursive substructures on top of ReAct-style agents, descending into sub-tasks to isolate local context in complex problems, but this recursion is primarily used for internal context management and the overall interaction pattern remains a standard step-by-step action loop.
 It is highlighted that the core limitation across all existing paradigms remains this rigid separation of plan and action.
@@
 We conduct experiments across three text-simulated environments that represent diverse decision-making challenges for LLM-based agents.
-ALFWorld (Shridhar et al., 2021) provides an embodied household environment where agents navigate and manipulate objects to complete daily tasks.
-WebShop (Yao et al., 2022) simulates an online shopping scenario requiring agents to search, compare, and purchase products based on specific requirements.
-ScienceWorld (Wang et al., 2022) presents a scientific laboratory setting where agents conduct experiments and manipulate scientific instruments.
+ALFWorld (shridhar2021alfworld) provides an embodied household environment where agents navigate and manipulate objects to complete daily tasks.
+WebShop (yao2022webshop) simulates an online shopping scenario requiring agents to search, compare, and purchase products based on specific requirements.
+ScienceWorld (wang2022scienceworld) presents a scientific laboratory setting where agents conduct experiments and manipulate scientific instruments.
 All three environments are partially observable Markov decision processes where agents cannot fully infer the environment state from observations alone.
@@
 To evaluate the effectiveness of ReCode, we divide our experiments into the inference part and the training part.
-For inference experiments, we compare against several mainstream paradigms: ReAct (Yao et al., 2023), which alternates between reasoning and action for iterative environment interaction; CodeAct (Wang et al., 2024b), extending ReAct’s action space from natural language to Python code.
-And some of the work focused on improving LLM-based agent planning: AdaPlanner (Sun et al., 2023), which pre-writes action sequences in Python and iteratively modifies them based on environmental feedback; and ADaPT (Prasad et al., 2024), which decomposes tasks on-demand using ReAct as the sub-task executor.
+For inference experiments, we compare against several mainstream paradigms: ReAct (yao2023react), which alternates between reasoning and action for iterative environment interaction; CodeAct (wang2024executable), extending ReAct’s action space from natural language to Python code.
+And some of the work focused on improving LLM-based agent planning: AdaPlanner (sun2023adaplanner), which pre-writes action sequences in Python and iteratively modifies them based on environmental feedback; and ADaPT (prasad2024adapt), which decomposes tasks on-demand using ReAct as the sub-task executor.
 | Method | Metric | ALFWorld | ScienceWorld | WebShop |
@@
 | Avg. Reward% (Std) | 100.0∗ | 88.5 14.58 | 97.5 7.05 |
-For training experiments, which evaluate how well a paradigm’s data structure lends itself to learning, we conduct supervised fine-tuning (SFT) (Wei et al., 2022) experiments on ReCode, ReAct, and CodeAct to directly compare their learning efficiency.
-Also, we compare some advanced training methods on the ReAct-style agent, including ETO (Song et al., 2024), which learns from expert trajectories via Direct Preference Optimization (Rafailov et al., 2023); and WKM (Qiao et al., 2024), introducing parameterized world models for training on expert and synthetic trajectories.
+For training experiments, which evaluate how well a paradigm’s data structure lends itself to learning, we conduct supervised fine-tuning (SFT) (weifinetuned) experiments on ReCode, ReAct, and CodeAct to directly compare their learning efficiency.
+Also, we compare some advanced training methods on the ReAct-style agent, including ETO (song2024trial), which learns from expert trajectories via Direct Preference Optimization (rafailov2023direct); and WKM (qiao2024agent), introducing parameterized world models for training on expert and synthetic trajectories.
 Due to differences in models, training setups, or environment configurations, we could not directly implement these latter works, but list them for reference.
@@
 Implementation Details
-For inference experiments, all methods use GPT-4o mini (Hurst et al., 2024) to ensure fair comparison.
+For inference experiments, all methods use GPT-4o mini (hurst2024gpt4o) to ensure fair comparison.
 We adapt few-shot prompts from prior work, making minimal modifications to fit each method’s format while keeping the number of examples consistent.
 The few-shot examples used for ReCode are provided in Appendix A.3.
-For training experiments, we use Qwen2.5-7B-Instruct (Yang et al., 2024a) as the base model.
-We construct the training datasets through the following process: First, for each training instance, we use a powerful teacher model, DeepSeek-V3.1 (DeepSeek-AI, 2024), to generate trajectories using ReAct, CodeAct, and ReCode respectively.
+For training experiments, we use Qwen2.5-7B-Instruct (yang2024qwen2) as the base model.
+We construct the training datasets through the following process: First, for each training instance, we use a powerful teacher model, DeepSeek-V3.1 (deepseekai2024deepseekv3technicalreport), to generate trajectories using ReAct, CodeAct, and ReCode respectively.
 Second, we filter these trajectories by keeping only the top 40% based on final reward.
@@
 ReCode generates significantly fewer but more compact data pairs (e.g., 6,385 pairs with 4.96M tokens on ALFWorld) compared to ReAct (27,607 pairs with 18.96M tokens).
-All models are trained using full-parameter fine-tuning on 8 NVIDIA A800 80GB GPUs with the open-source verl framework (Sheng et al., 2024).
+All models are trained using full-parameter fine-tuning on 8 NVIDIA A800 80GB GPUs with the open-source verl framework (sheng2024hybridflow).
 4.2 Inference Results
@@
 | CodeAct Gemini 2.5 Flash | 52.86 | 63.43 | 23.89 | 20.64 | 43.67 | 40.9 |
+| AdaPlanner Gemini 2.5 Flash | 77.86 | 85.82 | 30.41 | 33.25 | 8.68 | 47.2 |
+| ADaPT Gemini 2.5 Flash | 64.29 | 73.88 | 37.81 | 31.44 | 33.95 | 48.3 |
 | ReCode Gemini 2.5 Flash | 90.00 | 96.27 | 54.35 | 43.77 | 46.57 | 66.2 |
@@
 | CodeAct DeepSeek-V3.1 | 51.43 | 59.70 | 35.96 | 27.91 | 36.50 | 42.3 |
+| AdaPlanner DeepSeek-V3.1 | 87.86 | 92.54 | 27.75 | 26.25 | 36.65 | 54.2 |
+| blue ADaPTDeepSeek-V3.1 | 32.86 | 32.09 | 35.40 | 34.81 | 31.93 | 33.4 |
 | ReCode DeepSeek-V3.1 | 95.71 | 94.78 | 50.20 | 50.21 | 55.07 | 69.2 |
 Main Results
-As shown in Table 2, ReCode achieved significant performance improvements across all three environments, with an average score of 60.8, surpassing the best baseline method by 10.5 (relative 20.9%).
+As shown in Table 2, ReCode achieved significant performance improvements across all three environments, with an average score of 60.8, surpassing the best baseline method by approximately 10.5 (relative 20.9%).
 Notably, compared to the ReAct and CodeAct paradigms, ReCode improved average performance by 13.4 and 16.9, respectively.
@@
 On WebShop, ReCode maintains competitive performance with a score of 39.97, representing a 21.9% improvement over ADaPT.
-In the ScienceWorld environment, while ReCode shows comparable performance to ReAct on seen data and notable improvement on unseen tasks.
+In the ScienceWorld environment, while ReCode shows comparable performance to ReAct on seen data, it demonstrates notable improvement on unseen tasks.
 These consistent improvements across varied environments underscore the effectiveness of our approach in enhancing both task-specific performance and cross-domain adaptability.
-We provide detailed case studies in Section 4.4, showcasing representative trajectories from ALFWorld to illustrate ReCode’s decision-making process.
+We provide detailed case study in Appendix A.5, showcasing representative trajectory from ALFWorld to illustrate ReCode’s decision-making process.
 Multiple Model Results
-To illustrate that the ReCode paradigm is applicable to most models, we also experiment with Gemini 2.5 Flash (Comanici et al., 2025) and DeepSeek-V3.1, with results shown in Table 2.
-The cross-model evaluation demonstrates ReCode’s robustness and generalizability beyond a single LLM.
+To illustrate that the ReCode paradigm is applicable to most models, we also experiment with Gemini Flash 2.5 (comanici2025gemini) and DeepSeek-V3.1 (deepseekai2024deepseekv3technicalreport), with results shown in Table 2.
+The cross-model evaluation demonstrates ReCode’s robustness and generalizability beyond a single language model architecture.
 On Gemini 2.5 Flash, ReCode achieved an average performance of 66.2, maintaining consistent improvements over baseline methods.
-Similarly, when applied to DeepSeek-V3.1, it obtained a 69.2 average score, further validating the paradigm’s effectiveness.
-These results indicate that ReCode’s performance gains are not dependent on specific model, but rather stem from the fundamental improvements in the reasoning and code generation process.
+Similarly, when applied to DeepSeek-V3.1, our approach obtained a 69.2 average score, further validating the paradigm’s effectiveness.
+These results indicate that ReCode’s performance gains are not dependent on specific model characteristics or training procedures, but rather stem from the fundamental improvements in the reasoning and code generation process.
 Recursion Depth Analysis
@@
 Instead of the extensive, token-intensive, step-by-step reasoning cycles characteristic of reactive methods, ReCode’s hierarchical decomposition allows the agent to make more abstract and potent decisions.
-This leads to shorter, more direct reasoning paths, significantly reducing the number of API calls and total tokens required to solve a task while achieving superior performance.
+This leads to shorter, more direct reasoning paths, significantly reducing the total tokens required to solve a task while achieving superior performance.
 | Method | ALFWorld | ScienceWorld | WebShop | Average | ||
@@
 4.3 Training Results
-This section demonstrates the effectiveness of ReCode’s hierarchical data structure for agent training.
-We present comprehensive results across three environments and analyze the relationship between data and performance.
 | Method | ALFWorld | ScienceWorld | WebShop | Average | ||
@@
 On individual benchmarks, ReCode+SFT shows dominant performance on ALFWorld and WebShop, and remains competitive on ScienceWorld.
-Data Efficiency Analysis
-To systematically investigate the source of ReCode’s superior data efficiency, we conducted training on data subsets filtered by reward percentile .
-Figure 4 and Table 5 present the results on ScienceWorld.
-The results reveal distinct scaling behaviors.
-ReCode achieves 55.34% on seen tasks at using only 1,713 pairs, improving to 59.85% at with 3,500 pairs.
-In contrast, ReAct requires 12,833 pairs at to reach 60.28% on seen tasks.
-At matched percentiles, ReCode consistently outperforms ReAct with substantially fewer data pairs.
-At , ReCode achieves 44.87% with 688 pairs versus ReAct’s 34.05% with 3,094 pairs (4.5 more).
-At , ReCode reaches 55.34% with 1,713 pairs versus ReAct’s 36.34% with 6,320 pairs (3.7 more).
-Figure 4 shows that ReCode’s curves rise steeply then plateau, while ReAct’s curves show gradual improvement requiring substantially more data, confirming that ReCode’s hierarchical structure provides richer learning signals per training example.
 Figure 4: Performance of ReCode (blue) and ReAct (red) on ScienceWorld under different filtering percentiles.
@@
 Table 5: Training data statistics and corresponding ScienceWorld performance for ReAct and ReCode.
-The table lists the number of data pairs and the resulting average rewards for each method at different filtering percentiles ().
-4.4 Case Study
-To provide a concrete illustration of ReCode’s decision-making process, we analyze a representative trajectory from the ALFWorld environment for the task “put two alarmclock in dresser”.
-The complete execution trace and a diagram of the code generation flow are available in Figure 5 of the Appendix A.4.
-Initially, the agent formulates hybrid decisions that integrate abstract sub-task placeholder functions with concrete, primitive actions.
-These initial decisions, shown in the following code block, include functions like find_and_take for complex procedures that require further decomposition, and known, simple steps like run('go to dresser 1') as primitive actions.
-The agent then executes this hybrid code block sequentially.
-Primitive actions like run(...) are executed directly by the environment executor.
-While, placeholder functions such as find_and_take trigger an on-demand expansion, where the agent generates the necessary code to complete that specific sub-task just in time.
-This hybrid strategy allows the agent to maintain a high-level, strategic view of the task while committing to concrete actions only when the steps are simple and certain.
-It combines the robustness of hierarchical planning with the efficiency of direct execution, providing a flexible and powerful problem-solving framework that can adapt its level of abstraction as needed.
+The table lists the number of data pairs and the resulting average rewards (%) for each method at different filtering percentiles ().
+Data Efficiency Analysis
+To systematically investigate the source of ReCode’s superior data efficiency, we conducted training on data subsets filtered by reward percentile .
+Figure 4 and Table 5 present the results on ScienceWorld.
+The results reveal distinct scaling behaviors.
+ReCode achieves 55.34% on seen tasks at using only 1,713 pairs, improving to 59.85% at with 3,500 pairs.
+In contrast, ReAct requires 12,833 pairs at to reach 60.28% on seen tasks.
+At matched percentiles, ReCode consistently outperforms ReAct with substantially fewer data pairs.
+At , ReCode achieves 44.87% with 688 pairs versus ReAct’s 34.05% with 3,094 pairs (4.5 more).
+At , ReCode reaches 55.34% with 1,713 pairs versus ReAct’s 36.34% with 6,320 pairs (3.7 more).
+Figure 4 shows that ReCode’s curves rise steeply then plateau, while ReAct’s curves show gradual improvement requiring substantially more data, confirming that ReCode’s hierarchical structure provides richer learning signals per training example.
+4.4 Evaluation on Embodied Manipulation Tasks
+To further evaluate whether ReCode’s unified granularity control benefits tasks that require both high level planning and low level motor control, we conduct experiments on the ManiSkill (tao2025maniskill) benchmark.
+Unlike ALFWorld, WebShop, and ScienceWorld, ManiSkill requires continuous control of robot end effectors and fine grained pose adjustments, which provides a natural setting where dynamic adjustment of decision granularity is essential.
+We use two standard manipulation tasks, PickCube v1 and PullCube v1, based on the official ManiSkill3 Franka Panda robot.
+ReCode interacts with the environment through a small set of textual actions with continuous parameters; details of the wrappers, low level APIs, and task specifications are provided in Appendix A.1.
+| Model | Method | PickCube | PullCube |
+| GPT-4o mini | ReAct | 0 | 2 |
+| CaP | 34 | 34 | |
+| ReCode | 78 | 46 | |
+| Gemini 2.5 Flash | ReAct | 40 | 30 |
+| CaP | 82 | 44 | |
+| ReCode | 90 | 100 |
+We compare ReCode with two widely used paradigms for embodied decision making, ReAct and Code-as-Policies (CaP) (liang2023code), where ReAct performs step by step reasoning and emits actions, and CaP represents policies as full programs and recursively fills in missing functions before execution.
+We evaluate all methods with two LLMs, GPT-4o mini and Gemini 2.5 Flash, and report the success rate (%) on 50 episodes with random initialization under fixed seed for each task.
+The results in Table 6 show that ReAct struggles on ManiSkill, CaP improves performance but still relies on static programs that cannot adapt during execution, while ReCode consistently achieves the best results.
+The experiment indicate that ReCode’s unified recursive code based framework transfers to embodied manipulation and that flexible control of decision granularity yields practical benefits beyond text only environments.
 5 Conclusion
@@
 6 Limitations and Future Work
-While ReCode demonstrates significant potential, its effectiveness depends critically on the base model’s reasoning capabilities and few-shot example quality.
-The paradigm requires models to generate structured code within a specific framework, and any deficiency in planning ability compromises subsequent decision quality.
-As observed in our experiments, ReCode’s performance can be inconsistent, largely due to models’ difficulty adhering to the required code generation format or fully grasping the framework’s underlying logic.
-These limitations point to clear research directions centered on agent learning.
-A primary focus should be enhancing models’ intrinsic ability to understand and operate within the ReCode framework through specialized pre-training objectives or comprehensive fine-tuning strategies.
-Reinforcement learning approaches could reward agents for creating efficient hierarchical plans, enabling them to optimize their own expansion processes.
-Additionally, exploring methods to make the framework more robust to generation errors and investigating automatic curriculum learning for progressive skill acquisition represent promising avenues.
-These directions would represent significant steps toward more autonomous and capable agent learning systems.
+While ReCode shows strong potential, its effectiveness still depends heavily on the underlying language model and the quality of the few shot examples.
+The framework requires the model to produce structured code in a specific format, so weaknesses in planning ability or deviations from the required format can lead to unstable performance.
+A natural direction for future work is to strengthen the model’s understanding of the ReCode framework through targeted pre training or fine tuning, and to explore reinforcement learning signals that reward efficient hierarchical plans and reliable expansions.
+It is also promising to develop mechanisms that increase robustness to generation errors and to investigate curriculum style training procedures that gradually introduce more complex tasks and reasoning patterns.
 Appendix A Appendix
@@
 ALFWorld
-ALFWorld (Shridhar et al., 2021) provides a text-only game environment where the agent is required to complete a series of long-horizon household tasks based on natural language instructions.
-Built upon the embodied vision benchmark ALFRED (Shridhar et al., 2020), ALFWorld inherits its six core task types: pick and place, examine in light, clean and place, heat and place, cool and place, and pick two and place.
+ALFWorld (shridhar2021alfworld) provides a text-only game environment where the agent is required to complete a series of long-horizon household tasks based on natural language instructions.
+Built upon the embodied vision benchmark ALFRED (shridhar2020alfred), ALFWorld inherits its six core task types: pick and place, examine in light, clean and place, heat and place, cool and place, and pick two and place.
 After executing a command (e.g., “go to countertop 1”), the agent receives immediate textual feedback from the environment (e.g., “You arrive at countertop 1.
@@
 WebShop
-WebShop Yao et al.
-(2022) is an online shopping website environment constructed with 1.18 million real-world product data points.
+WebShop yao2022webshop is an online shopping website environment constructed with 1.18 million real-world product data points.
 The agent must navigate the site by searching and clicking buttons to browse various products, with the goal of finding and purchasing a specific item that meets the given requirements.
-We follow the data partitioning from MPO (Xiong et al., 2025), using a test set of 200 instances.
+We follow the data partitioning from MPO (xiong2025mpo), using a test set of 200 instances.
 WebShop provides a dense reward between 0 and 1 based on the agent’s actions and interactions with products.
@@
 The tasks in ScienceWorld are meticulously categorized, with each task having several variations and offering numerous different sub-goals.
-We adopt the data split from ETO (Song et al., 2024) and exclude Task-9 and Task-10 from the trajectories.
+We adopt the data split from ETO (song2024trial) and exclude Task-9 and Task-10 from the trajectories.
 ScienceWorld provides a dense reward between 0 and 1 for each task, based on the completion of its sub-goals.
@@
 | ScienceWorld | 1483 | 194 | 211 |
+ManiSkill
+We evaluate embodied decision making using two manipulation tasks from ManiSkill (tao2025maniskill), PickCube v1 and PullCube v1, both implemented with the official Franka Panda robot.
+Each environment exposes a continuous control interface based on incremental end–effector pose and gripper commands.
+To make this interface compatible with LLMs, we wrap the ManiSkill environments and provide a small set of textual control actions with continuous parameters.
+These actions allow the LLM to issue fine grained motor commands through executable code.
+We expose four textual actions that map directly to the underlying ManiSkill action vector: move(dx, dy, dz) specifies continuous translation of the end effector, rotate(d_roll, d_pitch, d_yaw) specifies continuous rotation increments, and gripper(opening) sets a continuous gripper opening value in .
+All parameters are real-valued floats and are clipped to the action bounds enforced by ManiSkill.
+This interface preserves the continuous nature of the control problem.
+Besides, to provide a compact textual observation to the LLM, we extract and serialize key environment states including the end–effector position, object position, goal position, and (for PickCube) the grasp status.
+This produces a consistent text description for each step while avoiding direct exposure of raw simulator states.
 A.2 Prompts
@@
 Here we present a representative chemistry task: creating green paint through color mixing.
-A.3.1 WebShop Examples
+WebShop Examples
 The WebShop examples demonstrate e-commerce navigation tasks where the agent must search for products, evaluate options against specified requirements, and complete purchases.
 Note that since clickable buttons in the environment are represented with square brackets (e.g., [Search], [Buy Now]), we use INPUT: and OUTPUT: instead of [Input] and [Output] in these examples to avoid confusion.
-A.4 Case Study
+A.4 Quantitative Analysis of ReCode
+To provide quantitative evidence for how ReCode adjusts its decision granularity during execution, we report statistics computed over trajectories on the unseen splits from ALFWorld, ScienceWorld and WebShop using GPT-4o mini.
+These metrics characterize the hierarchical structure that emerges from ReCode’s recursive expansion process.
+We measure three structural properties of the generated decision trees and these metrics capture the balance between coarse and fine decisions across tasks.
+Average recursion depth, defined as the maximum expansion depth reached in an episode, averaged over all trajectories;
+Average branching factor, defined as the average number of child functions or substeps produced by each expansion;
+Placeholder ratio, defined as the fraction of generated functions that remain as high-level placeholders rather than being expanded into primitive actions.
+| Environment | Avg. Depth () | Avg. Branching () | Placeholder Ratio |
+| ALFWorld | 0.19 | ||
+| WebShop | 0.17 | ||
+| ScienceWorld | 0.11 |
+Discussion
+The statistics reveal clear differences in how ReCode structures its decisions across environments.
+ALFWorld has the smallest average depth and branching factor, suggesting that its tasks rarely require deeply nested plans and instead emphasize structured exploration and local interaction in relatively simple scenes.
+WebShop uses deeper trees with larger branching factors, which is consistent with multi step product search, filtering, and comparison that naturally induce several layers of decision points.
+ScienceWorld exhibits the deepest and most complex decision structures, reflecting its multi stage experimental procedures and frequent decomposition into intermediate subgoals.
+Across all environments, the placeholder ratios remain relatively low and tend to decrease as task complexity increases.
+A plausible explanation is that more complex tasks require richer environment dynamics to support reliable planning, so ReCode executes more primitive actions and interacts more frequently with the environment before or alongside high level abstractions.
+These patterns provide quantitative evidence that ReCode adapts its decision structure and effective granularity to the structure and complexity of each environment.
+A.5 Case Study
+Initially, the agent formulates a hybrid plan that integrates abstract sub-tasks with concrete, primitive actions.
+This initial plan, shown in the following code block, includes distinct sub-procedures like find_and_take for complex procedures that require further decomposition, while directly embedding known, simple steps like run('go to dresser 1') as primitive actions within the main strategic flow.
+The agent then executes this hybrid plan sequentially.
+Primitive actions like run(...) are executed directly by the environment interpreter.
+In contrast, placeholder functions such as find_and_take trigger an on-demand expansion, where the agent generates the necessary code to complete that specific sub-task just in time.
+This hybrid strategy allows the agent to maintain a high-level, strategic view of the task while committing to concrete actions only when the steps are simple and certain.
+It combines the robustness of hierarchical planning with the efficiency of direct execution, providing a flexible and powerful problem-solving framework that can adapt its level of abstraction as needed.
 To provide a concrete illustration of ReCode’s decision-making process, we analyze a representative trajectory from the ALFWorld environment for the task “put two alarmclock in dresser”.